1. Your group will need approximately 6 g of the copper ammine complex to perform the analyses. The general formula for the copper ammine complex is [Cu(NH3)x]SO4•H2O Calculate the formula weight for several values of x, x=2 through x=6.

2. You will need to prepare about 6 g of the copper ammine product. Estimate how much CuSO4•5H2O you should start with to prepare approximately 6 g of the copper ammine complex.

3. Calculate how many grams of copper sulfate pentahydrate are required to prepare a 0.05 M solution in a 100-mL volumetric flask.

4. Show the equation needed to calculate the volume of 0.05 M stock solution needed to prepare a set of diluted solutions in a 10-mL volumetric flask.

5. Rearrange the equation from Pre-Lab Assignment 4 to show that you can quickly calculate the actual concentration on the basis of the actual volume of stock solution (0.05 M) delivered. (M1*V1 = M2V2)
Answer #1

The molar mass when X =2

  [Cu(NH3)2]SO4 H2O ,The molecular weight of complex = 63.55 + 2(17) +32.07 + 4(16) + (18) = 211 g/mol

b)The molecular weight when X = 3

   [Cu(NH3)3]SO4 H2O ,The molecular weight of complex = 63.55 + 3(17) +32.07 + 4(16) + (18) = 228.62 g/mol

c) The molecular weight when X = 4

  [Cu(NH3)4SO4.H2O ,The molecular weight of complex = 63.55 + 4(17) +32.07 + 4(16) +(18) =245.62 g/mol

d) The molecular weight when X = 5

  [Cu(NH3)5]]SO4 H2O ,The molecular weight of complex = 63.55 + 5(17) +32.07 + 4(16) + (18) = 262.62 g/mol

e) The molecular weight when X = 6

  [Cu(NH3)6]SO4 H2O ,The molecular weight of complex = 63.55 + 6(17) +32.07 + 4(16) + (18) = 279.62 g/mol

According to HOMEWORKLIB POLICY only 1 question per post is allowed. please post the rest question seperately.

PLEASE LEAVE THE POSITIVE FEEDBACK IF YOU LIKE THE SOLUTION